OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

put values in it and get result.. (y-y1)=m(x-x1) here m=slope

OpenStudy (anonymous):

your answer will be B...

OpenStudy (haseeb96):

(y-y1)=m(x-x1) just put the given slope and the point in it and get the answer